OpenStudy (anonymous):

A farmer has 1000 feet of fence to enclose a rectangular area. what dimensions for the rectangle result in the maximum area enclosed by the fence

OpenStudy (ranga):

Very much like the problem we just did. Assume the length of the rectangle to be L and the Width to be W. What is its perimeter in terms of L and W?

OpenStudy (anonymous):

(2L)(2W)

OpenStudy (anonymous):

sorry plus

OpenStudy (anonymous):

(2L)+(2W)

OpenStudy (ranga):

Yes. Equate that to 1000 and find W in terms of L.

OpenStudy (anonymous):

(L-1000)= W?

OpenStudy (ranga):

Perimeter = 2L + 2W = 2(L + W) = 1000 W = ?

OpenStudy (anonymous):

2L-1000=W?

OpenStudy (ranga):

First divide both sides of the equation by 2. Then solve for W.

OpenStudy (anonymous):

i have L+W=500 but i don't know how to get rid of the L but i have a feeling that i should divide the 500 by 2 and ill get 250

OpenStudy (ranga):

L + W = 500 We are not solving for L. We just want to find W in terms of L Subtract L from both sides of the equation and you will have W = ?

OpenStudy (anonymous):

so i would have 4W=250

OpenStudy (anonymous):

sorry 4w=100 but then i would have to divide by four on both sides

OpenStudy (ranga):

No. L + W = 500 Subtract L from both sides. W = 500 - L

OpenStudy (anonymous):

oh okayi see

OpenStudy (ranga):

So we have a rectangle whose Length = L Width = (500 - L) What is the area A of the rectangle?

OpenStudy (anonymous):

(L)(W)

OpenStudy (ranga):

Yes. Area = LW = L(500 - L) Follow?

OpenStudy (anonymous):

would i have to distribute?

OpenStudy (ranga):

Not necessary. We need to maximize the area which means we have to maximize the product L(500 - L) But you will notice the similarity to the previous problem where the product of two numbers x(30 -x) has to be maximized and the answer was x = 30/2 = 15. In the same way the product L(500 - L) will be maximum when L = 500/2 = 250. Follow?

OpenStudy (anonymous):

some what

OpenStudy (ranga):

Previously we plugged in various values of x from 1 to 17 and determined the maximum. But with a large number such as 500 it is going to be too time consuming without using calculus. So whenever a perimeter of a rectangle is fixed (as in this problem when perimeter = 1000) you will get maximum area when the length and the width of the rectangle are the same. In other words the area has to be a square. So if you put L = 250 in the equation L+W = 500 we had earlier you will find W is also 250. So Area is maximum when L = W = 250 Area = LW = (250)(250) = 62,500.

OpenStudy (anonymous):

oh okay well ill b

OpenStudy (anonymous):

racticing with oth

OpenStudy (ranga):

Alright.

OpenStudy (anonymous):

thank you

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."